OpenStudy (anonymous):

subtract and simplify if possible 5/11y^2-11y - 6/11y-11

jimthompson5910 (jim_thompson5910):

\[\Large \frac{5}{11y^2-11y} - \frac{6}{11y-11}\] \[\Large \frac{5}{y(11y-11)} - \frac{6}{11y-11}\] \[\Large \frac{5}{y(11y-11)} - \frac{6y}{y(11y-11)}\] \[\Large \frac{5-6y}{y(11y-11)}\] \[\Large \frac{5-6y}{11y^2-11y}\] So \[\Large \frac{5}{11y^2-11y} - \frac{6}{11y-11}=\frac{5-6y}{11y^2-11y}\]